User name Saha
Log entry time 16:05:43 on January26,2001
Entry number 55791
0800: Run 1750 continuing (Kin 2)
0915: Run 1750 ended. 368k events
MCC takes away beam for injector studies (~15 mins)
0935: Run 1751 started.
1042: Shift checklist completed (R. Gilman)
1100: Run 1751 ended. 269k events
MCC takes away beam to recalibrate the MPS BCM system(~1 hr)
Changing to BeO target tro check beam spot position
13:45 Beam back, looked ok on BeO (raster was still on though).
Asked MCC to let us move to LH2 target. They will call us back.
13:50 MCC called - we can move the target.
14:00 Target moved
Run 1752 started.
1513 Run 1752 ended, 500k events
1517 Run 1753 started
1600 Run 1753 continuing (~275k)